Find 5A + 3B.
A = 2 -5
-5 2
B = 5 -2
2 -5

Answer by Alan3354(69443)

Find 5A + 3B.
A =
+2 -5
-5 2
5A =
10 -25
-25 10
---------
B =
5 -2
2 -5
3B =
15 -6
6 -15
-----
5A + 3B =
+25 -31
-19 -5